OpenStudy (anonymous):

gilbert scored half of his team points.juan scored 8 fewer points than gilbert. the team scored 36 points how many poitns did juan score

OpenStudy (anonymous):

Gilbert = 36/2 = 18 Juan = 18 - 8 = 10 So Juan scored 10 points

OpenStudy (anonymous):

10